OpenStudy (anonymous):

the average of x, y and z is one-third greater than x. the sum y+z is what percent greater than x?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {x + y + z \over 3} = {1 \over 3} + x }\) \(\Large \color{Black}{\Rightarrow 3x + 1 = x + y + z }\) \(\Large \color{Black}{\Rightarrow 2x + 1 = y + z }\) Hmm

OpenStudy (pfenn1):

I think you meant \[{x + y + z \over 3} =(1+{1 \over 3}) x \]

OpenStudy (anonymous):

the sum y+z is what percent greater than x?

OpenStudy (anonymous):

???? Percent???

OpenStudy (anonymous):

Till here I figured out?!!! I cudn't figure out what percent????

OpenStudy (apoorvk):

Just a sec - confirming - Is Parth's version the right one or is pfenn's? #"one third greater than x"

OpenStudy (pfenn1):

Simplify the equation I gave you and then rearrange so that you have y+z on the left side of the equation and see what you come out with.

OpenStudy (anonymous):

@pfenn1 Simplified and i got y+z = 3x that means the sum of y and z is equal to 3 times x or 300% of x. Am i right?

OpenStudy (pfenn1):

Correct!

OpenStudy (anonymous):

One thing I can't figure out: When it's written greater than do we have to add?

OpenStudy (anonymous):

?

OpenStudy (pfenn1):

Mostly, but it really depends on the wording. The problem that @ParthKohli had was that he wasn't multiplying x by the fraction and that you have to add the fraction to 1 because the average was 1/3 great than x. If you just multiply by 1/3, you will get an answer that is 1/3 of x but you really want 1/3 more than x. So you have to multiply by 1+(1/3).

OpenStudy (anonymous):

that means the sum is one third of x greater than x?

Parth (parthkohli):

So that can be said as 4/3x right?

OpenStudy (anonymous):

@pfenn1 please reply @ParthKohli yeah

Parth (parthkohli):

Then do it :P

OpenStudy (anonymous):

did it told the answer already above.

OpenStudy (pfenn1):

Hi. Sorry, I had an appointment. Yes, Zeerak